OpenStudy (anonymous):

use the substitution method to solve the system of equations. write your answer as an ordered pair 2) x+2y=2 y=2x+6

OpenStudy (anonymous):

WHy hello there. Here's how you use the substitution method: 1) Pick your favorite equation of the two and your favorite variable. 2) Solve the equation for that variable. 3) Substitute into the OTHER equation, which gives you an equation with just 1 variable in it. 4) Solve for that variable =D 5) Use the variable you know, substitute into either of the first two equations and solve for the other variable.

OpenStudy (anonymous):

how do i find the variable tho?

OpenStudy (anonymous):

By followin' my steps =) Let me see your attempt.

OpenStudy (anonymous):

okay..... x+2y=2 (-2)+2(2)=2 (2)=2(-2)+6 :D

OpenStudy (anonymous):

i need help with this one........... x-2y=3 x+y=1